Time Remaining 55 minutes 7 seconds

00:55:07

Item 17

Item 17

Time Remaining 55 minutes 7 seconds

00:55:07

Tulsa Company, (a merchandising Co.) has the following data pertaining to the year ended December 31, 2019: (CPA adapted)
  

Purchases $ 620,000
Beginning inventory 204,000
Ending inventory 227,000
Freight-in 67,000
Freight-out 83,500

  
What is the cost of goods sold for the year?

2. Luxus, Inc. employs 42 sales personnel to market its line of luxury automobiles. The average car sells for $28,000, and a 6 percent commission is paid to the salesperson. Luxus, Inc. is considering a change to the commission arrangement where the company would pay each salesperson a salary of $2,700 per month plus a commission of 2% of the sales made by that salesperson. The amount of total monthly car sales at which Luxus, Inc. would be indifferent as to which plan to select is:

3. Dorcan Corporation manufactures and sells T-shirts imprinted with college names and slogans. Last year, the shirts sold for $12.60 each, and the variable cost to manufacture them was $7.60 per unit. The company needed to sell 23,400 shirts to break-even. The after tax net income last year was $6,060. Donnelly's expectations for the coming year include the following: (CMA adapted)

  • The sales price of the T-shirts will be $15.
  • Variable cost to manufacture will increase by one-third.
  • Fixed costs will increase by 10%.
  • The income tax rate of 40% will be unchanged.


Sales for the coming year are expected to exceed last year's by 1,170 units. If this occurs, Dorcan's sales volume in the coming year will be:
Answer #1
Ans 1.
Tulsa Company
a Beginning Inventory                         204,000
b Add : Purchase                         620,000
c Add: Freight In                           67,000
d Less : Ending Inventory                         227,000
Cost of Goods Sold for the Year =a+b+c-d=                         664,000
Ans 2.
Assume Indifferent car sales per month =x
Monthly Sales Commission in existing system =x*28000*6%
Monthly sales commission by proposed system
=42*2700+x*28000*2%
At Point of indifference both commission will be same.
So , x*28000*6%=42*2700+x*28000*2%
x*1680=113400+560*x
x=101.25
So Indifferent no of Cars sales /month=101.25=101 rounded
Ans 3.
Finding Fixed Operating cost /year
a Unit sales Price $                         12.60
b Less Unit Variable cost $                           7.60
c Unit Contribution Margin $                           5.00
d BEP no of T Shirts = $                 23,400.00
e annual Fixed cost=c*d= $               117,000.00
Let us Find Last Year Sales Volume
f After Tax Net Income $                   6,060.00
g Tax Rate = 40%
h EBT =f/60%= $                 10,100.00
i Annual Fixed cost $               117,000.00
j Annual Contribution Margin =h+i= $               127,100.00
c Unit Contribution Margin $                           5.00
l Units sold last year =j/c=                     25,420.00 units
m Coming year sale will exceed last yr sale by                        1,170.00 units
n Coming Year sales Volume =l+m=                     26,590.00 units